OpenStudy (anonymous):

divide ^3sqrt243a^7b^5 / ^3sqrt9a^5b^4

OpenStudy (turingtest):

Dear lord, please learn to use the equation feature, this is very hard to decipher and hurts my eyes to look at.

OpenStudy (turingtest):

Is this your problem?\[\sqrt[3]{243}a^7b^5 \over \sqrt[3]9a^5b^4\]

OpenStudy (gg):

how to write equation in question?

OpenStudy (anonymous):

yses

OpenStudy (anonymous):

\[3a ^{2/3}b ^{1/3}\]

OpenStudy (turingtest):

\[{\sqrt[3]{243}a^7b^5\over \sqrt[3]{9}a^5b^4}=\sqrt[3]{243\over9}a^{7-5}b^{5-4}=\sqrt[3]{27}a^2b=3a^2b\]

OpenStudy (turingtest):

@Gg I don't know how to write an equation into the question, I would just post it after.